OpenStudy (anonymous):

Water is pumped out of a holding tank at a rate of 5-5e^-0.09t liters/minute, where t is in minutes since the pump is started. If the holding tank contains 1000 liters of water when the pump is started, how much water does it hold one hour later? Round your answer to one decimal place. Amount of water in tank after one hour ? liters.

OpenStudy (anonymous):

can u help me please

OpenStudy (istim):

What is E?

OpenStudy (anonymous):

its something that can be put on a calculator. Idk.

OpenStudy (anonymous):

its on a calculator

OpenStudy (istim):

I don't think I can help much. My opinion is to substitute 60 for t, then divide 1000 by that. But that's obviously incorrect.

OpenStudy (anonymous):

hello coco are u gonna anser?

OpenStudy (anonymous):

hello

OpenStudy (istim):

Give him/her some time.

OpenStudy (istim):

They have a long answer for you. Patience may help.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

I have got the method already and I am now solving the equation

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

volume of the water pumped out =integral 5-5e^-0.09t = 5t - 5(-1/0.09) e^(-0.09t) after 1 hour which is 60min , volume pumped out= 5(60)-5(-1/0.09)e^((-0.09)(60)) = 300.3 volume remained = 1000-300.3= 699.7

OpenStudy (anonymous):

sorry that it spends you quite a long time to wait. I thought in wrong way at the beginning but now it should be correct.

OpenStudy (anonymous):

is 699.7 the answer. I ahve to get it right

OpenStudy (anonymous):

Chlorophyll help meee someone!

OpenStudy (anonymous):

OMG can one of u just say something please.

OpenStudy (anonymous):

lol is anyone uh there? I have four problems left on this quiz that i hae been doing for like almost four hours and its only ten questions! im now irritated

OpenStudy (anonymous):

I hate u, the answer was wrong. FML

OpenStudy (anonymous):

wrong?

OpenStudy (anonymous):

= 5∫ ( 1 - e ^ (-.09t ) dt [ 0, 60 ] = 5 ( t + e ^ (-.09t ) / .09 ) [ 0, 60 ] = 244.7 L Volume remain after 1 hr: 1,000 - 244.7 = 755.3

OpenStudy (anonymous):

Hey, we do the best, but not guarantee it's 100% correct. It's your resposibility to double check.